OpenStudy (kymber):

\(f(27)=\) \(\sqrt{2x-5}\)

OpenStudy (ajprincess):

What do u need to find?

OpenStudy (arlen):

put x= 27

OpenStudy (kymber):

Find f(27) it says.. I get to \(16=\)\(\sqrt{x}\) And don't know what to do then.

OpenStudy (arlen):

f(27)=7

OpenStudy (kymber):

How, nya?

OpenStudy (arlen):

\[f(27)= \sqrt{2*27-5}= \sqrt 49=7\]

OpenStudy (kymber):

Oh. You put 27 in for x? :I

OpenStudy (kymber):

Thank you.

OpenStudy (arlen):

yes

OpenStudy (arlen):

:)

OpenStudy (ajprincess):